OpenStudy (anonymous):

The perimeter of a rectangular store receipt is ten and five-sixths inches. If three of the sides total eight and one-half inches, what is the length of the fourth side? two and one-sixth inches two and one-third inches 3 inches three and one-twelfth inches

OpenStudy (jonnyvonny):

You know the perimeter of a rectangle is: P=2l+2w. However, in this case, since they give us 3 of the 4 sides, are enabled to assume that P=x+3sides; P=x+8.5. They give us P...10(5/6)=x+8(1/2)....from there, just solve for x. If you need help with getting a common denominator, feel free to reply. :)